OpenStudy (gabylovesyou):

A company's monthly profit, P, from a product is given by P = −x2 + 105x − 1050, where x is the price of the product in dollars. What is the lowest price of the product, in dollars, that gives a monthly profit of $1,550? (Do not put a $ in your answer.) @jim_thompson5910

OpenStudy (gabylovesyou):

is it 40 ?

OpenStudy (compassionate):

I don't know, Gaby, is it? Show us your work.

OpenStudy (gabylovesyou):

1550=-x^2+105x-1050 x^2-105x+2600=0 (x-40)(x-65) = 0 x = 40

OpenStudy (compassionate):

Subtract 1,550 x cannot be negative so multiply it by -1 Factor Solve for x Did I read that right?

OpenStudy (gabylovesyou):

am i wrong lol

OpenStudy (compassionate):

Nope. (;

OpenStudy (gabylovesyou):

yay me

OpenStudy (agent0smith):

@Gabylovesyou i wouldn't have given you a medal if you were wrong :P http://www.wolframalpha.com/input/?i=1550%3D-x%5E2%2B105x-1050

OpenStudy (gabylovesyou):

:DDD